OpenStudy (anonymous):

write each equation in the system of equations in slope intercept form. use the slope intercept form to determine the number of solutions the system has. A. no solution. B. one solutions. C. infinite number of solutions. D, cannot be determined.

OpenStudy (anonymous):

6x-4y=9

OpenStudy (anonymous):

and -18x+12y=-18

OpenStudy (openstudier):

Slope-intercept form: \[y=mx+b\]

OpenStudy (openstudier):

We can put both equations in terms of y in order to get slope-intercept form!

OpenStudy (anonymous):

y=6/4x+9?

OpenStudy (anonymous):

y=3/4x-18

OpenStudy (openstudier):

Nope, try again.

OpenStudy (anonymous):

try getting rid of the coefficient of the y first.

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

divide the whole equation by the coefficient of y.

OpenStudy (anonymous):

is it no solution?